An object is 0.25 m from a concave mirror with a focal length of 0.6 m the object is 0.38 m talk.


1
Expert's answer
2021-03-08T08:09:56-0500

Given,

Object distance (u) = 0.25m

Focal length of the concave mirror "(f)=0.6m"

Length of the object "(h_o)=0.38m"

Let position of image be at v distance from the concave mirror.

Now, applying the mirror formula,

"\frac{1}{v}+\frac{1}{u}=\frac{1}{f}"

Now, substituting the values,

"\frac{1}{v}=\frac{1}{f}-\frac{1}{u}"

for the concave mirror, focal length always will be negative. so

"\frac{1}{v}=\frac{1}{0.6}-\frac{1}{0.25}"


"\Rightarrow \frac{1}{v}=\frac{0.25-0.6}{0.150}"


"\Rightarrow \frac{1}{v}=\frac{-0.35}{0.15}"

"v=\frac{-3}{7}cm"

Hence the height of the image "h_i=\frac{-h_o\times v}{u}=\frac{3\times 0.38}{7}cm"

"h_i=\frac{1.14}{7}cm =0.16cm"